OpenStudy (anonymous):

Find the equation of a line with the given points. (5, 4) (5, 6)

OpenStudy (aravindg):

well use two point form

OpenStudy (aravindg):

\[y-y_1= \frac{(x-x_1)(y_2-y_1)}{x_2-x_1}\]

OpenStudy (hba):

Well you can do it this way or in a two step method. 1) Find slope \[m=\frac{ y_2-y_1 }{ x_2-x_1 }\] 2)put in the pont slope form \[y-y_1=m(x-x_1)\] The other method as shown by @aravindg is the two point form \[y-y_1=\frac{ y_2-y_1 }{ x_2-x_1 }(x-x_1)\]

OpenStudy (hba):

Here (x1,y1)=(5,4) (x2,y2)=(5,6)

OpenStudy (aravindg):

well two point form is the direct approach when you are given two points ! : )